Fermi National Laboratory


Accelerator Update for February 23, 2005 - February 25, 2005

Wednesday February 23
The day shift began with Operations monitoring store 4000 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and Meson MCenter experiments taking beam.

Operations reported at 10:24 AM that the Neutron Therapy Facility (NTF) began taking beam.

Tevatron (TeV) experts began an end-of-store study at 10:39 AM.

Operation terminated store 4000 at 11:04 AM.

TeV damper studies began at 11:21 AM.

Operations loaded the TeV at 12:16 PM with a 36x0 study-store.

TeV studies ended at 12:51 PM. Cryo system technicians began their work on the TeV sector B1 wet engine.

Operations prepared the Collider Detector at Fermilab (CDF) for an access at 12:55 PM.

Operations put the TeV through a dry squeeze at 4:38 PM.

Operations put the TeV into shot setup at 5:03 PM. Shot setup halts Switchyard (SY) beam.

Operations established store 4002 at 7:05 PM with an initial luminosity of 107.43E30. This set a NEW LUMINOSITY RECORD. Pbar resumed stacking.

Operations resumed sending beam to MCenter at 7:29 PM.

At 8:41 PM, the Main Injector (MI) lost its beam permit due to MECAR. This occurred four more times over the next hour.

At 10:39 PM, Operations reported that MECAR had caused several more MI trips.

At 11:15 PM, Operations rebooted MECAR, but then had trouble recovering from the reboot. The investigation continued.

At 11:57 PM, operators power cycled and rebooted MECAR locally. Operators contacted experts.

Thursday February 24
The midnight shift began with Operations monitoring store 4002, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

An MI expert began investigating the MECAR problem around 1:45 AM. He left about an hour and a half later without discovering the cause.

Operations put the Recycler into shot setup, to stash antiprotons, at 3:44 AM. The stash was completed by 4:37 AM and Pbar resumed stacking.

Operations reestablished beam to MCenter at 5:05 AM.

At 5:27 AM, Operations reported that they'd been having problems keeping MiniBooNE running since establishing beam to SY. Operators disabled the SY beam to investigate and discovered a quadrupole power supply (QPS229) not regulating properly. Operators replaced a control card.

MiniBooNE began taking beam at 6:13 AM. MCenter resumed taking beam at 6:19 AM.

Operations established beam to MTest at 10:02 AM.

NTF requested the use of Linac beam at 10:15 AM.

At 1:44 PM, Operations turned off the Linac beam to allow work on a quadrupole power supply and on the Booster RF low-level system. The Linac work was completed within fifteen minutes.

Booster completed their RF work at 2:22 PM. Pbar resumed stacking and Operations resumed sending beam to MiniBooNE, MTest, and MCenter.

At 3:16 PM, MiniBooNE began tripping off every couple of pulses. Operators lowered the experiment's intensity to investigate.

At 3:55 PM, operators asked for help from a MiniBooNE expert to diagnose the problem.

Experimenters at MTest reported at 7:04 PM that they were through with the beam for the night.

Operations terminated store 4002 at 8:15 PM. Operators made the D-Zero (D0) collision hall ready for an access.

MCenter experimenters reported at 9:04 PM that they would be down for a few hours. Operators disabled the SY beam.

D0 came out of access and Operations put the TeV through a dry squeeze at 11:18 PM.

Operations put the TeV into shot setup at 11:44 PM.

Friday February 25
The midnight shift began with the TeV in shot setup, with Operations monitoring the stack and stash, and with MiniBooNE taking beam.

At 12:29 AM, Operations reported that MiniBooNE would not stay on. Operators continued their investigation.

Operations established store 4004 at 2:06 AM with an initial luminosity of 83.6E30. The lower luminosity was due to the loss of two antiproton transfers.

At 2:31 AM, Operations reported that they were still having problems with MiniBooNE. They couldn't get a hold of a MiniBooNE expert and so asked a Booster expert for help.

Operations established beam to MCenter at 3:19 AM.

At 4:06 AM, Operations halted Linac beam while the Booster expert replaced a quadrupole power supply (Q803) in the MI-8 transfer line.

Stacking and MiniBooNE beam resumed at 5:50 AM.

The Plan for Friday and the Weekend
The plan for today is to terminate the store this evening and then let D0 make an access. The rest of the weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/25/2005   email Fermilab